OpenStudy (anonymous):

The price of a suit decreased from $280 to $182 during a clearance sale. What is the percent of decrease in the price? Round to the nearest percent, if necessary. A. 98% B. 65% C. 54% D. 35% plzz help

OpenStudy (anonymous):

its D! :)

OpenStudy (anonymous):

@geny55 Thx^_^

OpenStudy (anonymous):

medal?

OpenStudy (anonymous):

do you need me to explain my answer? @music.1.drug

OpenStudy (anonymous):

The price decreased by $98. The original price was $280 so it is the percentage decrease of this figure we are looking for. The calculation is:- 98/280 x 100 = 9800/280 = 35 so the decrease is 35%

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

i explained it :)